A 15-story (equal size floors) office building has 10% nonrentable space, and the building lobby (151,200 square feet, rentable) occupies the first three floors. Company LYNX is interested in leasing one of the six equal sized usable areas for tenants on the sixth floor [floor total usable area = rentable square feet of the floor – common space of the floor]. The sixth floor’s common space is 22% of the floor’s rentable square feet. The leasing broker indicated that LYNX’s annual rent will be $199,500. The owner will apply load factor for the common rentable space on the sixth floor as well as the load factor for the building’s lobby to calculate the rentable area for LYNX. (a) What is the Gross Area of the Building? (b) What is the annual base rent per square feet of rentable space, rounded to zero decimal places, for LYNX?
Answer #1

Answer:

Part a:

gross area of the building = 2,520,000 square feet

.

Part b:

annual base rent per square feet = $0.97

annual base rent per square feet = $1 (rounded to zero decimal places)

Explanation

sixth floor total rentable space = 151200

common space = 22%*151200 = 33264

total usable area = 151200-33264 = 117936

load factor = rentable area/usable area = 151200/117936 = 1.28

.

Hence rentable area = 1.28*151200 = 193846.20

annual base rent per square feet will be 193846.20/199500 = $0.97

annual base rent per square feet = $1 (rounded to zero decimal places)